OpenStudy (anonymous):

Find P (rolling 2 or 5) with one number cube. A.1/2 B.1/3 C. /6 D. 1

OpenStudy (ganpat):

is it a regular cube ?? 1 to 6 ??

OpenStudy (anonymous):

i not sure

OpenStudy (ganpat):

ok, in that case, (if that is true) Probability will be 1/6 + 1/6 = 2/6 = 1/3... 2nd option.. 1/6 = probability of getting 2 or 5 from a cube of 6.. and since its or.. it has been added...

OpenStudy (amistre64):

while i agree that that is the correct procedure; it really didnt give the asker a chance to participate and work thru it to develop the concepts better in their own mind ....